# Multi-Agent Large Language Models for Zero-Shot Robotics Manipulation

### v1

@@ IV-D Results for Zero-Shot Evaluation
Table I presents results achieved by the three baseline methods, along with our proposed MALMM, across 9 different tasks. From the table, we observe that MALMM outperforms all baselines across all 9 tasks, including long-horizon tasks such as stack blocks and empty container, as well as tasks involving complex shapes, such as ’meat off grill’ and ’rubbish in bin’. The Code as Policies is able to generate a successful trajectory for only two tasks. This limited success is because the original Code as Policies implementation relied on few-shot examples to perform well on tasks involving regularly shaped objects. In our evaluation, we replaced these few-shot examples with coordinate definitions and detailed instructions about the functions available for the LLM to call. However, Code as Policies completely fails in this zero-shot setting, which was also reported in [17].
VoxPoser [34], which generates 3D voxel maps for value functions to predict waypoints, successfully generates trajectories for three tasks with good accuracy. For two of these tasks, it was the second-best performing method. However, its performance did not generalize well to other tasks.
Both the single-agent (SA) and our proposed multi-agent framework, MALMM, successfully generate trajectories for all 9 tasks. However, MALMM consistently outperforms the SA approach by using different agents for specific roles, enabling it to plan and to execute long-horizon tasks effectively while detecting and correcting intermediate failures. We provide qualitative comparison between Single Agent and MALMM for the Stack Blocks task in Appendix-D.
@@ IV-E Multi-Agent Ablation
We perform ablation for each of the components (agents) in MALMM to evaluate how each of them contributes to the overall performance. We consider two tasks, namely stack blocks and empty container and report results in Table II.
We first analyze the importance of intermediate environment feedback. To this end, we consider the Single Agent (SA) setting and remove the environmental feedback provided after each intermediate step. In this setup LLM generates the full manipulation plan at once and executes it without revisions. As shown in Table II, the Single Agent without environmental feedback exhibits 12% and 24% drop in performance for the ’stack blocks’ and ’empty container’ tasks respectively. By analyzing failure cases of both methods, we observed that the environmental feedback provided after each intermediate step crucially effects the agent’s ability to detect unforeseen situations and to recover from failures such as collisions and inaccurate grasping.
@@
| Single Agent (SA) | ✗ | ✗ | ✗ | ✗ | 8.00 | 12.00 |
| Single Agent (SA) | ✓ | ✗ | ✗ | ✗ | 20.00 | 36.00 |
| Multi-Agent (MA) | ✓ | ✓ | ✓ | ✗ | 36.00 | 48.00 |
| MALMM | ✓ | ✓ | ✓ | ✓ | 56.00 | 64.00 |
We next validate the advantage of the Multi-Agent architecture with separate LLM agents for planning and code generation. As shown in Table II, the Multi-Agent system (MA), consisting of a dedicated Planner and Coder, demonstrates 16% and 12% performance improvement over SA for the two tasks respectively. This can be attributed to the inherent limitations of LLMs in managing very long context conversations [16]. In the SA setup, where a single LLM is responsible for both high-level planning and low-level code generation, the agent must handle an extensive context, particularly for long-horizon tasks. This often leads to errors such as failing to account for collisions with other objects, omitting input arguments for predefined functions, using variables before they are initialized, and even forgetting the specified goal. In contrast, the MA system mitigates these issues by dividing the workload among specialized agents. The Planner and Coder agents in the MA setting focus on specific roles through specialized prompts and communicate with each other reducing the likelihood of errors and hallucinations, in particular for longer tasks.
Our initial Multi-Agent system pre-defines the cyclic sequence of the Planner, the Coder, and the Coder Executor, see Fig. 3(b), assuming each agent correctly completes its task. However, hallucinations may occur even within multi-agent systems [15]. For example, the Coder agent may miss the variable initialization resulting in compilation errors or incomplete sub-goal code generation, such as producing the code only for approaching the object without grasping it. In such situations, the Coder may need to be re-executed to correct for errors before passing control to the Planner. To automate this process, we introduce a Supervisor agent that dynamically re-routs the execution process to the next agent based on the input instruction, entire communication history of all active agents, and the role descriptions of all agents, rather than following a fixed sequence. This adaptive approach at the core of our MALMM method improves the performance of the dual-agent MA setup by 20% in the ’stack blocks’ task and 16% in the ’empty container’ task respectively, see Table II.
@@ IV-F MALMM is Better at Long-Horizon Planning
To validate the effectiveness of MALMM in long-horizon tasks, we create three variations of the ’stack blocks’ task, each with a different number of blocks, and compare the performance of MALMM to the Single Agent setup. Results in Fig. 5 indicate that while the Single Agent setup struggles with stacking 3 and 4 blocks, MALMM significantly outperforms SA, in particular for tasks that require longer planning.
@@ IV-G Results for Vision-Based Observations
To approach real-world settings, where direct access to the environment states is not available, we next perform experiments restricted to vision-based observations in the form of 3D point clouds. We evaluate the performance of MALMM and compare it to the Single Agent setup on three tasks: ’put block’, ’rubbish in bin’, and ’close jar’. Consistently with our previous state-based experiments, results of MALMM in Table III significantly outperform the Single-Agent baseline across all three tasks and confirm the advantage of our proposed Multi-Agent framework.
By comparing results in Tables I and III we observe degradation of performance when switching to vision-based observations. This can be attributed to inaccuracies of vision-based estimators such as 3D bounding box detection and grasp estimation. Note that our current vision pipeline makes use of single-view scene observations. A parallel work, Manipulate Anything [57] showed that a relatively straightforward extension to multi-view settings can reduce the impact of occlusions and result in higher accuracy.
@@
| Single Agent | 24.00 | 68.00 | 40.00 |
| MALMM | 56.00 | 84.00 | 52.00 |
@@ V-A Limitations
Despite its significant advantages over a Single LLM Agent, MALMM has several limitations. First, MALMM relies on three gpt-4-turbo agents, making it costly to operate. Using open-source LLMs such as LLaMA [58] could help reduce costs, but running these models locally also incurs expenses. Second, like other LLM-based planners, MALMM depends on manual prompt engineering, which impacts its performance. However, advancements in alignment [59] and prompting [13, 14] reduce reduced efforts. Finally, MALMM requires accurate bounding box estimation to determine correct grasp positions, but as the complexity of objects increases, bounding boxes alone may not provide enough information for precise grasping. Our experiments with vision pipeline suggest that using pretrained grasping and placement models, such as M2T2 [52], could improve performance for complex manipulation tasks.
V-B Conclusions
This work explored LLM agents for solving previously unseen manipulation tasks. In particular, we have proposed the first multi-agent LLM framework for robotics manipulation MALMM and demonstrated its advantages over the single agent baselines. Our method uses task-agnostic prompts and requires no in-context learning examples for solving new tasks. Extensive evaluation demonstrated excellent results of MALMM for a variety of manipulation tasks, especially for tasks that require longer horizon planning. Future work will consider tasks with richer object interactions and will extend our experiments to real robotics scenes.
@@ V-B Conclusions
The Supervisor prompt of MALMM, shown in Figure 7, contains the same four variables as the Single Agent prompt described in Appendix-C1: [INSERT TASK], [INSERT EE POSITION], [INSERT EE ORIENTATION], and [STATE]. Additionally, the prompts for Planner and Coder in MALMM are depicted in Figures 8 and 9 respectively.
@@
We evaluated on nine RLBench tasks which are listed in Table IV along with the task instruction and success criteria.
@@
To obtain the visual observations, we use RGBD frames from the front view. First, we generate 3D point clouds from the depth channel of input RGBD frames. Then, we segment 3D point clouds of each source object by projecting LangSAM [51] generated 2D masks into 3D. Next, we employ M2T2 [52] model to identify potential object grasp poses. We filter M2T2 grasp poses to retain the ones with a top-down gripper orientation. Finally, we apply DBSCAN [60] clustering to the numerous grasp poses and select the central pose. For the placement, we use 3D bounding box of the target object to identify the placement area. Thus, the selected grasp pose of the object to be moved, along with the 3D bounding box of the target object are used as the environment state. Figure 11 illustrates the process of selecting the environment state, as described for the close jar task with the instruction close red jar with lid. In addition, to dynamically update the grasp and placement states, we introduce a history-aware mechanism to mitigate jitter in the extracted visual states (i.e., grasp and placement poses). Specifically, during each execution, if the translation distance was less than 0.01 meters or the change in the gripper orientation was below 30 degrees, we retain the visual state from the previous step. This approach helps maintaining stability during manipulation.

### v2

@@ IV-D Results for Zero-Shot Evaluation
Table I presents the results for the three baseline methods, along with our proposed MALMM, across 9 different tasks. From the table, we observe that MALMM outperforms all baselines across all 9 tasks, including long-horizon tasks such as stack blocks and empty container, as well as tasks involving complex shapes, such as meat off grill and rubbish in bin. Moreover, the Code as Policies is able to generate a successful trajectory for only two tasks. This limited success is because the original Code as Policies implementation relied on few-shot examples to perform well on tasks involving regularly shaped objects. In our evaluation, we replaced these few-shot examples with coordinate definitions and detailed instructions about the functions available for the LLM to call. However, Code as Policies completely failed in this zero-shot setting, which was also reported in [11].
VoxPoser [16], which generates 3D voxel maps for value functions to predict waypoints, successfully generated trajectories for three tasks with good accuracy. For two of these tasks, it was the second-best performing method. However, its performance did not generalize well to other tasks.
Both the SA and our proposed multi-agent framework, MALMM, successfully generated trajectories for all 9 tasks. However, MALMM consistently outperformed the SA approach by using different agents for specific roles, enabling it to generate accurate high-level plans and low-level code while mitigating hallucinations. In addition to the gpt-4-turbo experiments, we evaluated MALMM using the open-source LLaMA-3.3-70B [29]; the results are presented in Table I. Although there is a drop in performance compared to gpt-4-turbo, MALMM, with LLaMA-3.3-70B, outperforms the existing baselines by a sizable margin. We provide qualitative comparison between Single Agent and MALMM for the Stack Blocks task in Appendix-D.
@@ IV-E Multi-Agent Ablation
We performed ablation for each of the components (agents) in MALMM in order to evaluate how each of them contributes to the overall performance. We considered two tasks, namely stack blocks and empty container, and report the results in Table II.
We first analyzed the importance of the intermediate environment feedback. To this end, we considered the Single Agent (SA) setting and removed the environment feedback provided after each intermediate step. In this setup, LLM generates the full manipulation plan at once and executes it without revisions. As shown in Table II, the SA without environment feedback exhibits 12% and 24% drop in performance for the stack blocks and empty container tasks, respectively. By analyzing the failure cases of both methods, we observed that the environment feedback provided after each intermediate step crucially affected the agent’s ability to detect unforeseen situations and recover from failures such as collisions and inaccurate grasping.
@@
| Single Agent (SA) | ✗ | ✗ | ✗ | ✗ | 0.08 | 0.12 |
| Single Agent (SA) | ✓ | ✗ | ✗ | ✗ | 0.20 | 0.36 |
| Multi-Agent (MA) | ✓ | ✓ | ✓ | ✗ | 0.36 | 0.48 |
| MALMM | ✓ | ✓ | ✓ | ✓ | 0.56 | 0.64 |
We next validated the advantage of the Multi-Agent architecture with separate LLM agents for planning and code generation. As shown in Table II, the MA, consisting of a dedicated Planner and Coder, demonstrated 16% and 12% performance improvement over SA for the two tasks respectively. This can be attributed to the inherent limitations of LLMs in managing very long context conversations [6]. In the SA setup, where a single LLM is responsible for both the high-level planning and the low-level code generation, the agent must handle an extensive context, particularly for long-horizon tasks. This often leads to errors such as failing to account for collisions with other objects, omitting the input arguments for the predefined functions, using variables before they were initialized, and even forgetting the specified goal. In contrast, the MA system mitigates these issues by dividing the workload among specialized agents. The Planner and the Coder agents in the MA setting focus on specific roles through specialized prompts and communicate with each other, thus reducing the likelihood of errors and hallucinations, in particular for longer tasks.
Our initial Multi-Agent system pre-defines the cyclic sequence of the Planner, the Coder, and the Coder Executor, see Fig. 3(b), assuming that each agent correctly completes its task. However, hallucinations may occur even within multi-agent systems [33]. For example, the Coder may miss the variable initialization resulting in compilation errors or incomplete sub-goal code generation, such as producing code only for approaching the object without grasping it. In such situations, the Coder may need to be re-executed in order to correct possible errors before passing the control to the Planner. To automate this process, we introduced a Supervisor agent that dynamically re-routes the execution process to the next agent based on the input instruction, the entire communication history of all active agents, and the role descriptions of all agents rather than following a fixed sequence. This adaptive approach is at the core of our MALMM framework, and it improves the performance of the dual-agent MA setup by 20% in the ’stack blocks’ task and 16% in the ’empty container’ task, respectively, as shown in Table II.
@@ IV-F MALMM is Better at Long-Horizon Planning
To validate the effectiveness of MALMM in long-horizon tasks, we created three variations of the ’stack blocks’ task, each with a different number of blocks, and compared the performance of MALMM to the Single Agent setup. The results in Fig. 5 indicate that while the Single Agent setup struggles with stacking 3 and 4 blocks, MALMM substantially outperforms SA, in particular for tasks that require longer planning. Fig. 1 illustrates an example of failure recovery with key steps by MALMM compared to SA on the stack block task.
@@ IV-G Results for Vision-Based Observations
In order to approach real-world settings where direct access to the environment states is not available, we next perform experiments in simulation restricted to vision-based observations in the form of 3D point clouds. We evaluate the performance of MALMM and compare it to the Single Agent setup on three tasks: (ii) put block, (ii) rubbish in bin, and (iii) close jar. Consistently with our previous state-based experiments, the results in MALMM in Table III show sizable improvements over the Single-Agent baseline across all three tasks and confirm the advantage of our proposed Multi-Agent framework.
By comparing the results in Tables I and III we observe degradation in performance when switching to vision-based observations. This can be attributed to inaccuracies of the vision-based estimators such as 3D bounding box detection and grasp estimation. Note that our current vision pipeline makes use of single-view scene observations. A parallel work, Manipulate Anything [34] showed that a relatively straightforward extension to multi-view settings can reduce the impact of occlusions and yield higher accuracy.
@@
| Single Agent [11] | 0.24 | 0.68 | 0.40 |
| MALMM | 0.56 | 0.84 | 0.52 |
IV-H Results for the Real-World Experiments
For the real-world robotics setup, we evaluated five tasks as shown in Fig. 6: close jar (put lid on top of jar), put block (place a block in the red target area), put rubbish in bin (place rubbish in the bin), put case (place an earbuds case in the red target area), and put jar in bin (place a jar in the bin) – in ten different initial states, each with both MALMM and the Single Agent. As shown in Table IV, consistently with our simulation results, MALMM outperforms the Single Agent on all five tasks by a sizable margin. It achieved a 40% higher success rate for both put case and rubbish in bin, 30% higher success rate for put block, and 20% higher success rate for close jar and jar in bin. To further demonstrate zero-shot capabilities of our method, Fig. 7 demonstrates successful performance of MALMM for three new tasks, each defined by high-level user instruction.
| Agents | Close | Put | Rubbish | Put | Jar |
| Jar | Block | in Bin | Case | in Bin | |
| Single Agent [11] | 2/10 | 3/10 | 2/10 | 3/10 | 3/10 |
| MALMM | 4/10 | 6/10 | 6/10 | 7/10 | 5/10 |
@@ V-A Limitations
Despite its string advantages over SA, MALMM has several limitations. First, MALMM relies on three gpt-4-turbo agents, making it costly to operate. Using open-source LLMs is possible at the cost of reduced accuracy (cf. Table I). Second, like other LLM-based planners, MALMM depends on manual prompt engineering, which impacts its performance. However, advancements in prompting [5] can reduce these efforts. Finally, MALMM requires accurate bounding box estimation to determine the correct grasp positions, but as the complexity of the objects increases, the 3D bounding boxes alone may not provide enough information for precise grasping. Our experiments with a vision pipeline in simulation and in a real-world scenario suggest that using pretrained grasping and placement models, such as M2T2 [27], could improve the performance for complex manipulation tasks. Even though MALMM performs over 6 better than the SA on the long-horizon task (stacking 4 blocks), it achieves only a 28% success rate, revealing limitations in handling rich object interactions. Additionally, the vision-based pipeline is not robust to severe occlusions due to its single-view limitation.
V-B Conclusion and Future Work
We explored the use of LLM agents for solving previously unseen manipulation tasks. In particular, we proposed the first multi-agent LLM framework for robotics manipulation MALMM and demonstrated its advantages over single-agent baselines. Our method uses task-agnostic prompts and requires no in-context learning examples for solving new tasks. Extensive evaluations, both in simulation and real-world settings, demonstrated excellent results for MALMM for a variety of manipulation tasks. In future work, we will explore richer object interactions and more complex tasks, including articulated objects. We also aim to use vision-language foundation models to incorporate contextual scene details beyond object bounding boxes.
@@ V-B Conclusion and Future Work
The Supervisor prompt of MALMM, shown in Figure 9, contains the same four variables as the Single Agent prompt described in Appendix-C1: [INSERT TASK], [INSERT EE POSITION], [INSERT EE ORIENTATION], and [STATE]. Additionally, the prompts for Planner and Coder in MALMM are depicted in Figures 10 and 11 respectively.
@@
We evaluated on nine RLBench tasks which are listed in Table V along with the task instruction and success criteria.
@@
To obtain the visual observations, we use RGBD frames from the front view. First, we apply gpt-4-turbo [28] on the RGB image and task instruction, prompting it to generate appropriate names for the source and target objects, which is helpful for LangSAM to segment. Second, we generate 3D point clouds from the input RGBD frames. Then, we segment the 3D point clouds of the source and target objects by projecting the 2D masks generated by LangSAM [26] into 3D. Next, we employ the M2T2 [27] model to identify potential grasp poses on the source objects. We filter these grasp poses to retain only those with a top-down gripper orientation. Finally, we apply DBSCAN [35] clustering to the numerous grasp poses and select the central pose. For the placement, we use the 3D bounding box of the target object to identify the placement area. Thus, the selected grasp poses of the objects to be moved, along with the 3D bounding box of the target object, are used as the environment state. Figure 13 illustrates the process of selecting the environment state, as described for the close jar task with the instruction close red jar with lid. In addition, to dynamically update the grasp and placement states, we introduce a history-aware mechanism to mitigate jitter in the extracted visual states (i.e., grasp and placement poses). Specifically, during each execution, if the translation distance is less than 0.01 meters or the change in gripper orientation is below 30 degrees, we retain the visual state from the previous step. This approach helps maintain stability during manipulation.